# Flavored QCD axion and Modular invariance

### v1

@@ IV.2 Neutrino mass and mixing
$\alpha_{e}=[0.7038,0.7042]\,,\quad\alpha_{e2}=[1.2247,1.3095]\,,\quad\alpha_{\mu}=[0.5917,0.5925]\,$,
$\alpha_{\tau}=[1.5475,1.5477]\,,\quad\alpha_{\tau 1}=[1.1786,1.1814]\,,\quad\alpha_{\tau 2}=[0.3896,0.4309]\,$,
$\arg(\alpha_{e})=[5.80,2\pi]\,,\quad\arg(\alpha_{e2})=[4.80,5.36]\,,\quad\arg(\alpha_{\mu})=[0.09,0.42]\,$,
$\arg(\alpha_{\tau 1})=[0.85,1.44]\,,\quad\arg(\alpha_{\tau 2})=[1.30,2.00]\,.$ (105)
@@
The seesaw mechanism in Eq.(96) operates at the $U(1)_{X}$ symmetry breakdown scale, while its implications are measured by experiments below the electroweak scale. Therefore, quantum corrections to neutrino masses and mixing angles can be crucial, especially for degenerate neutrino masses Antusch:2005gp . However, based on our numerical calculation that the neutrino mass spectra exhibit normal mass hierarchy (NH) at the scale of $U(1)_{X}$ breakdown (as depicted in the left panel of Fig.3), we can safely assume that the renormalization group running effect on observables can be ignored. To show the viable parameter space, we scan the precision constraints $\{\theta_{13}$, $\theta_{23}$, $\theta_{12}$, $\Delta m^{2}_{\rm Sol}$, $\Delta m^{2}_{\rm Atm}\}$ at $3\sigma$ from Table-3. Using the reference values from Eqs.(41) and (94) and taking $\langle\chi\rangle=10^{12}$ GeV of Eq.(92), see also Eq.(97), we determine the input parameter spaces in Eqs.(77) and (81) at the $U(1)_{X}$ breaking scale :
$\beta_{1e}=[0.479,0.578],\quad\tilde{\beta}_{1e}=[0.657,0.816],\quad\beta_{2\mu}=[1.115,1.311],\quad\tilde{\beta}_{2\mu}=[1.349,1.613],$
$\beta_{3\tau}=[1.149,1.415],\quad\tilde{\beta}_{3\tau}=[0.871,1.125],\quad\beta_{1\mu}=[0.977,1.132],\quad\tilde{\beta}_{1\mu}=[1.121,1.354],$
$\beta_{1\tau}=[0.460,0.656],\quad\tilde{\beta}_{1\tau}=[1.200,1.552],\quad\beta_{2e}=[1.036,1.352],\quad\tilde{\beta}_{2e}=[1.031,1.322]\,$,
$\beta_{2\tau}=[1.136,1.345],\quad\tilde{\beta}_{2\tau}=[1.311,1.540],\quad\beta_{3e}=[0.972,1.262],\quad\tilde{\beta}_{3e}=[0.394,0.645],$
$\beta_{3\mu}=[1.176,1.465],\quad\tilde{\beta}_{3\mu}=[0.472,0.669],$
$\arg(\beta_{1e})=[4.1,4.6],\quad\arg(\beta_{2\mu})=[\pi,3.7],\quad\arg(\beta_{3\tau})=[4.9,5.6],$
$\arg(\beta_{1\mu})=[1.4,2.3],\quad\arg(\beta_{1\tau})=[5.2,5.7],\quad\arg(\beta_{2e})=[2.5,\pi],$
$\arg(\beta_{2\tau})=[0,0.4]\cup[5.7,2\pi],\quad\arg(\beta_{3e})=[0,1.5],\quad\arg(\beta_{3\mu})=[5.5,2\pi],$ (106)
@@
$\gamma_{11}=[0.450,0.613],\quad\tilde{\gamma}_{11}=[0.657,0.810],\quad\gamma_{12}=[1.376,1.524],\quad\tilde{\gamma}_{12}=[0.390,0.614],$
$\gamma_{13}=[0.686,0.864],\quad\tilde{\gamma}_{13}=[1.055,1.294],\quad\gamma_{22}=[0.952,1.309],\quad\tilde{\gamma}_{22}=[0.938,1.160],$
$\gamma_{23}=[0.975,1.217],\quad\tilde{\gamma}_{23}=[1.341,1.555],\quad\gamma_{33}=[1.000,1.234],\quad\tilde{\gamma}_{33}=[1.151,1.352],$
$\arg(\gamma_{11})=[4.4,5.0],\quad\arg(\gamma_{12})=[\pi,3.8],\quad\arg(\gamma_{13})=[0,1.4],$
$\arg(\gamma_{22})=[1.8,2.6],\quad\arg(\gamma_{23})=[5.1,5.8],\quad\arg(\gamma_{33})=[2.1,2.8],$ (107)
@@
Neutrino oscillation experiments currently aim to make precise measurements of the Dirac CP-violating phase $\delta_{CP}$ and atmospheric mixing angle $\theta_{23}$. For the parameter regions explored in our model, referring to the two-dimensional allowed regions at $3\sigma$ presented in Ref.Esteban:2020cvm , we note that there are no favored regions with respect to $\theta_{23}$ and $\delta_{CP}$, as shown in the right panel of Fig.3. Ongoing experiments like DUNE DUNE:2018tke , as well as proposed next-generation experiments such as Hyper-K Hyper-Kamiokande:2018ofw , are poised to greatly reduce uncertainties in the values of $\theta_{23}$ and $\delta_{CP}$, providing appropriate data set for our proposed model. As shown in the left panel of Fig.4, the predicted $0\nu\beta\beta$-decay rate, plotted against $\delta_{CP}$, lies entirely below the experimental upper bound in Eq.(104). Moreover, ongoing and future experiments on $0\nu\beta\beta$-decay like NEXT NEXT:2020amj , SNO++SNO:2022trz , KamLAND-Zen KamLAND-Zen:2022tow , Theia Theia:2019non , SuperNEMO Arnold:2004xq may reach a sensitivity to test the NH of our model. Cosmological and astrophysical measurements provide powerful constraints on the sum of neutrino masses. The upper bound on the sum of the three active neutrino masses can be summarized as $\sum m_{\nu}=m_{\nu_{1}}+m_{\nu_{2}}+m_{\nu_{3}}<0.120$ eV at $95\%$ CL for TT, TE, EE+lowE+lensing+BAO Planck:2018vyg . The right panel of Fig.4 illustrates that the sum of neutrino masses lies in the range of $0.0634$ to $0.0756$ eV when $0.0036\lesssim m_{\rm lightest}=m_{\nu_{1}}\lesssim 0.0107$ eV.
@@
A phenomenologically viable model has been presented, in which $U(1)_{X}$ charges and modular weights $k_{I}$ are systematically assigned to all quark and lepton fields. We have shown that, unlike in ordinary cases, flavor-violating interactions of the flavored QCD axion to $s,d$-quarks and $\mu,e$-leptons are suppressed to $\mathcal{O}(\lambda^{4})$. This is a consequence of the $U(1)_{X}$ charge assignments and the properties of the fermion mixing matrices. This suppression makes rare decays $B^{\pm}\to K^{\pm}+a$ and $\tau\to\mu+a$ the dominant processes constraining the axion decay constant $f_{A}$, while kaon decays ($K^{+}\to\pi^{+}+a$) remain subdominant. So, the strongest experimental bounds in our model arise from the axion-electron coupling, which is tightly constrained by red giant branch star cooling observations.
Numerically, the model is consistent with current experimental data on flavor mixing and mass spectra. It successfully accounts for the observed quark Dirac CP phase and predicts a neutrino mass spectrum with a normal hierarchy. Future reductions in the uncertainties of atmospheric mixing angle $\theta_{23}$ and lepton Dirac CP phase $\delta_{CP}$ will provide appropriate data set for our proposed model. Additionally, it provides specific predictions for a flavored QCD axion, including a mass $m_{a}=1.82\times 10^{-4}$ eV and a photon coupling strength $|g_{a\gamma\gamma}|=1.75\times 10^{-14}~\text{GeV}^{-1}$, which are consistent with all known flavor constraints.
@@ Acknowledgements.
I would like to thank Prof. Eung Jin Chun (KIAS) for the thoughtful and in-depth discussions. This work is supported by NSFC 12135006.
@@ Appendix E Quark and charged-lepton mixing matrices
$V^{\ell}_{L}=\tilde{C}_{\ell}{\left(\begin{array}[]{ccc}1-\frac{1}{2}\theta^{2}_{2}&-\theta_{3}\,e^{i\varphi_{\ell}}&-\theta_{2}+\frac{1}{6}\theta^{3}_{2}\\ \theta_{3}\,e^{-i\varphi_{\ell}}&1-\frac{1}{2}\theta^{2}_{1}&-\theta_{1}\\ \theta_{2}-\frac{\theta^{3}_{2}}{6}&\theta_{1}(1-\frac{1}{2}\theta^{2}_{2})&1-\frac{1}{2}\theta^{2}_{2}\end{array}\right)}\tilde{K}_{\ell}+{\cal O}(\lambda^{4})\,$, (133)
where $\varphi_{\ell}=2\alpha^{\ell}_{3}-\alpha^{\ell}_{2}$, $\tilde{C}_{\ell}={\rm diag}(e^{i(\alpha^{\ell}_{2}-\alpha^{\ell}_{1}-\alpha^{\ell}_{3})},e^{i(\alpha^{\ell}_{3}-\alpha^{\ell}_{1})},e^{i(\alpha^{\ell}_{2}-\alpha^{\ell}_{1})})$, and $\tilde{K}_{\ell}={\rm diag}(e^{i(\alpha^{\ell}_{1}-2\alpha^{\ell}_{2})},1,e^{i2\alpha^{\ell}_{1}})$ with $\alpha^{\ell}_{1}=\frac{1}{2}\arg(\alpha^{\ast}_{\tau 2}Y^{(4)\ast}_{\bf 1})$, $\alpha^{\ell}_{2}\simeq\frac{1}{2}\arg(\alpha^{\ast}_{\tau 1}Y^{(4)\ast}_{\bf 1})-\frac{1}{2}\alpha^{\ell}_{1}$, $\alpha^{\ell}_{3}\simeq-\frac{1}{2}\arg(\alpha^{\ast}_{\tau 1}\alpha_{\tau 2}Y^{(4)\ast}_{\bf 1}Y^{(8)}_{\bf 1})-\frac{1}{2}\alpha^{\ell}_{1}-\frac{1}{2}\alpha^{\ell}_{2}$, and
$\theta_{1}\simeq\frac{|\alpha_{\tau 2}Y^{(8)}_{\bf 1}|}{|\alpha_{\tau}|}(2\,{\rm Im}\,\tau)^{-4}\,$,
$\theta_{2}\simeq\frac{|\alpha_{\tau 1}Y^{(4)}_{\bf 1}|}{|\alpha_{\tau}|}(2\,{\rm Im}\,\tau)^{-2}\sim\lambda\,$,
$\theta_{3}\simeq\frac{1}{2}\big|\frac{\alpha_{\tau 2}}{\alpha_{\tau}}Y^{(8)}_{\bf 1}\big|\frac{|\alpha^{\ast}_{\tau 1}\alpha_{\tau 2}Y^{(4)\ast}_{\bf 1}Y^{(8)}_{\bf 1}|}{|\alpha_{\mu}|^{2}}(2\,{\rm Im}\,\tau)^{-14}\Delta^{-8}_{\chi}\,.$ (134)
Here $\theta_{1}\sim{\cal O}(\lambda^{2})$, $\theta_{2}\sim{\cal O}(\lambda)$, and $\theta_{3}\sim{\cal O}(\lambda^{3-4})$ are roughly expected for $\Delta_{\chi}=0.634$ and ${\rm Im}\,\tau\simeq 1$. The right-handed charged-lepton mixing matrix $V_{R}^{\ell}$ reads
$V^{\ell}_{R}=\tilde{C}^{r}_{\ell}{\left(\begin{array}[]{ccc}1&0&0\\ 0&1-\frac{1}{2}\theta^{r2}_{23}&-\theta^{r}_{23}\\ 0&\theta^{r}_{23}&1-\frac{1}{2}\theta^{r2}_{23}\end{array}\right)}\tilde{K}^{r}_{\ell}+{\cal O}(\lambda^{4})\,$, (138)
where $\tilde{C}^{r}_{\ell}={\rm diag}(e^{-i(\tilde{\alpha}^{\ell}_{2}+\tilde{\alpha}^{\ell}_{3})},e^{i\tilde{\alpha}^{\ell}_{3}},e^{i\tilde{\alpha}^{\ell}_{2}})$, $\tilde{K}^{r}_{\ell}={\rm diag}(1,e^{-i\tilde{\alpha}^{\ell}_{1}},e^{i\tilde{\alpha}^{\ell}_{1}})$
$\theta^{r}_{23}=\frac{|\alpha_{\mu}\alpha^{\ast}_{\tau 2}Y^{(8)\ast}_{\bf 1}|}{|\alpha_{\tau}|^{2}}\Delta^{4}_{\chi}(2{\rm Im}\,\tau)^{-4}\sim{\cal O}(\lambda^{2})$ (139)
and, $\tilde{\alpha}^{\ell}_{1}=\frac{1}{2}\arg(\alpha^{\ast}_{\tau 2}\alpha_{\mu}Y^{(8)\ast}_{\bf 1})$, $\tilde{\alpha}^{\ell}_{2}\simeq\frac{1}{2}\arg(\alpha^{\ast}_{\tau 1}\alpha_{e}Y^{(4)\ast}_{\bf 1})-\frac{1}{2}\tilde{\alpha}^{\ell}_{1}$, $\tilde{\alpha}^{\ell}_{3}\simeq\frac{1}{2}\arg(\alpha^{\ast}_{\mu}\alpha_{e2}Y^{(4)}_{\bf 1})+\frac{1}{2}\tilde{\alpha}^{\ell}_{1}-\frac{1}{2}\tilde{\alpha}^{\ell}_{2}$.

### v2

@@ IV.2 Neutrino mass and mixing
$\alpha_{e}=[1.1477,1.1486]\,,\quad\alpha_{e2}=[1.0532,1.0633]\,,\quad\alpha_{\mu}=[0.6091,0.6095]\,$,
$\alpha_{\tau}=[0.3703,0.3707]\,,\quad\alpha_{\tau 1}=[0.3954,0.3975]\,,\quad\alpha_{\tau 2}=[0.7390,0.7419]\,$,
$\arg(\alpha_{e})=[1.26,1.46]\,,\quad\arg(\alpha_{e2})=[0.44,0.95]\,,\quad\arg(\alpha_{\mu})=[0.71,1.65]\,$,
$\arg(\alpha_{\tau 1})=[4.90,5.97]\,,\quad\arg(\alpha_{\tau 2})=[3.15,4.13]\,.$ (105)
@@
The seesaw mechanism in Eq.(96) operates at the $U(1)_{X}$ symmetry breakdown scale, while its implications are measured by experiments below the electroweak scale. Therefore, quantum corrections to neutrino masses and mixing angles can be crucial, especially for degenerate neutrino masses Antusch:2005gp . However, based on our numerical calculation that the neutrino mass spectra exhibit normal mass hierarchy (NH) at the scale of $U(1)_{X}$ breakdown (as depicted in the left panel of Fig.3), we can safely assume that the renormalization group running effect on observables can be ignored. To show the viable parameter space, we scan the precision constraints $\{\theta_{13}$, $\theta_{23}$, $\theta_{12}$, $\Delta m^{2}_{\rm Sol}$, $\Delta m^{2}_{\rm Atm}\}$ at $3\sigma$ from Table-3. Using the reference values from Eqs.(41,94) and taking $\langle\chi\rangle=2\times 10^{10}$ GeV of Eq.(92), see also Eq.(97), we determine the input parameter spaces in Eqs.(77) and (81) at the $U(1)_{X}$ breaking scale :
$\beta_{1e}=[0.642,0.668],\quad\tilde{\beta}_{1e}=[1.000,1.129],\quad\beta_{2\mu}=[1.470,1.479],\quad\tilde{\beta}_{2\mu}=[0.514,0.535],$
$\beta_{3\tau}=[0.998,1.122],\quad\tilde{\beta}_{3\tau}=[0.335,0.374],\quad\beta_{1\mu}=[1.428,1.464],\quad\tilde{\beta}_{1\mu}=[0.629,0.666],$
$\beta_{1\tau}=[1.184,1.230],\quad\tilde{\beta}_{1\tau}=[1.530,1.566],\quad\beta_{2e}=[1.204,1.230],\quad\tilde{\beta}_{2e}=[0.723,0.759]\,$,
$\beta_{2\tau}=[1.394,1.416],\quad\tilde{\beta}_{2\tau}=[0.671,0.697],\quad\beta_{3e}=[0.658,0.696],\quad\tilde{\beta}_{3e}=[0.946,1.135],$
$\beta_{3\mu}=[1.340,1.386],\quad\tilde{\beta}_{3\mu}=[1.274,1.338],$
$\arg(\beta_{1e})=[1.20,1.71],\quad\arg(\beta_{2\mu})=[5.50,2\pi],\quad\arg(\beta_{3\tau})=[0.69,1.17],$
$\arg(\beta_{1\mu})=[2.61,3.75],\quad\arg(\beta_{1\tau})=[2.60,3.20],\quad\arg(\beta_{2e})=[3.46,4.20],$
$\arg(\beta_{2\tau})=[5.57,2\pi],\quad\arg(\beta_{3e})=[5.71,2\pi],\quad\arg(\beta_{3\mu})=[2.50,3.56],$ (106)
@@
$\gamma_{11}=[0.616,0.659],\quad\tilde{\gamma}_{11}=[0.876,0.924],\quad\gamma_{12}=[1.331,1.379],\quad\tilde{\gamma}_{12}=[0.946,0.967],$
$\gamma_{13}=[0.768,0.828],\quad\tilde{\gamma}_{13}=[0.760,0.816],\quad\gamma_{22}=[1.426,1.468],\quad\tilde{\gamma}_{22}=[0.652,0.715],$
$\gamma_{23}=[1.365,1.409],\quad\tilde{\gamma}_{23}=[0.930,0.966],\quad\gamma_{33}=[1.343,1.413],\quad\tilde{\gamma}_{33}=[0.800,0.874],$
$\arg(\gamma_{11})=[1.70,2.40],\quad\arg(\gamma_{12})=[4.41,5.11],\quad\arg(\gamma_{13})=[3.70,4.40],$
$\arg(\gamma_{22})=[4.10,4.99],\quad\arg(\gamma_{23})=[5.34,6.10],\quad\arg(\gamma_{33})=[0.18,0.81],$ (107)
@@
Neutrino oscillation experiments currently aim to make precise measurements of the Dirac CP-violating phase $\delta_{CP}$ and atmospheric mixing angle $\theta_{23}$. For the parameter regions explored in our model, referring to the two-dimensional allowed regions at $3\sigma$ presented in Ref.Esteban:2020cvm , we note that there are no favored regions with respect to $\theta_{23}$ and $\delta_{CP}$, as shown in the right panel of Fig.3. Ongoing experiments like DUNE DUNE:2018tke , as well as proposed next-generation experiments such as Hyper-K Hyper-Kamiokande:2018ofw , are poised to greatly reduce uncertainties in the values of $\theta_{23}$ and $\delta_{CP}$, providing appropriate data set for our proposed model. As shown in the left panel of Fig.4, the predicted $0\nu\beta\beta$-decay rate, plotted against $\delta_{CP}$, lies entirely below the experimental upper bound in Eq.(104). Moreover, ongoing and future experiments on $0\nu\beta\beta$-decay like NEXT NEXT:2020amj , SNO++SNO:2022trz , KamLAND-Zen KamLAND-Zen:2022tow , Theia Theia:2019non , SuperNEMO Arnold:2004xq may reach a sensitivity to test the NH of our model. Cosmological and astrophysical measurements provide powerful constraints on the sum of neutrino masses. The upper bound on the sum of the three active neutrino masses can be summarized as $\sum m_{\nu}=m_{\nu_{1}}+m_{\nu_{2}}+m_{\nu_{3}}<0.120$ eV at $95\%$ CL for TT, TE, EE+lowE+lensing+BAO Planck:2018vyg . The right panel of Fig.4 illustrates that the sum of neutrino masses lies in the range of $0.0634$ to $0.0756$ eV when $0.0034\lesssim m_{\rm lightest}=m_{\nu_{1}}\lesssim 0.0065$ eV.
@@
A phenomenologically viable model has been presented, in which $U(1)_{X}$ charges and modular weights $k_{I}$ are systematically assigned to all quark and lepton fields. We have shown that, unlike in ordinary cases, flavor-violating interactions of the flavored QCD axion to $s,d$-quarks and $\mu,e$-leptons are suppressed to $\mathcal{O}(\lambda^{4})$. This suppression makes rare decay $B^{\pm}\to K^{\pm}+a$ the dominant processes constraining the axion decay constant $f_{A}$, while kaon decays ($K^{+}\to\pi^{+}+a$) remain subdominant. So, the strongest experimental bounds in our model arise from the axion-electron coupling, which is tightly constrained by red giant branch star cooling observations.
Numerically, the model is consistent with current experimental data on flavor mixing and mass spectra. It successfully accounts for the observed quark Dirac CP phase and predicts a neutrino mass spectrum with a normal hierarchy. Future reductions in the uncertainties of atmospheric mixing angle $\theta_{23}$ and lepton Dirac CP phase $\delta_{CP}$ will provide appropriate data set for our proposed model. Additionally, it provides specific predictions for a flavored QCD axion, including a mass $m_{a}=9.12\times 10^{-3}$ eV and a photon coupling strength $|g_{a\gamma\gamma}|=1.69\times 10^{-13}~\text{GeV}^{-1}$, which are consistent with all known flavor constraints.
@@ Acknowledgements.
I would like to thank Prof. Eung Jin Chun (KIAS) for the thoughtful and in-depth discussions, and Chris Ahn (Port moody secondary school) for pointing out some typos. This work is supported by NSFC 12135006.
@@ Appendix E Quark and charged-lepton mixing matrices
$V^{\ell}_{L}=\tilde{C}_{\ell}{\left(\begin{array}[]{ccc}1-\frac{1}{2}\theta^{2}_{2}&0&-\theta_{2}\\ 0&1-\frac{1}{2}\theta^{2}_{1}&-\theta_{1}\\ \theta_{2}&\theta_{1}(1-\frac{1}{2}\theta^{2}_{2})&1-\frac{1}{2}\theta^{2}_{2}\end{array}\right)}\tilde{K}_{\ell}+{\cal O}(\lambda^{4})\,$, (133)
where $\varphi_{\ell}=2\alpha^{\ell}_{3}-\alpha^{\ell}_{2}$, $\tilde{C}_{\ell}={\rm diag}(e^{i(\alpha^{\ell}_{2}-\alpha^{\ell}_{1}-\alpha^{\ell}_{3})},e^{i(\alpha^{\ell}_{3}-\alpha^{\ell}_{1})},e^{i(\alpha^{\ell}_{2}-\alpha^{\ell}_{1})})$, and $\tilde{K}_{\ell}={\rm diag}(e^{i(\alpha^{\ell}_{1}-2\alpha^{\ell}_{2})},1,e^{i2\alpha^{\ell}_{1}})$ with $\alpha^{\ell}_{1}=\frac{1}{2}\arg(\alpha^{\ast}_{\tau 2}Y^{(8)\ast}_{\bf 1})$, $\alpha^{\ell}_{2}\simeq\frac{1}{2}\arg(\alpha^{\ast}_{\tau 1}Y^{(4)\ast}_{\bf 1})-\frac{1}{2}\alpha^{\ell}_{1}$, $\alpha^{\ell}_{3}\simeq\frac{1}{2}\arg(\alpha^{\ast}_{e}\alpha_{e2}Y^{(4)}_{\bf 1})+\frac{1}{2}\alpha^{\ell}_{1}-\frac{1}{2}\alpha^{\ell}_{2}$, and
$\theta_{1}\simeq\frac{|\alpha_{\tau 2}Y^{(8)}_{\bf 1}|}{|\alpha_{\tau}|}(2\,{\rm Im}\,\tau)^{-4}\,,\theta_{2}\simeq\frac{|\alpha_{\tau 1}Y^{(4)}_{\bf 1}|}{|\alpha_{\tau}|}(2\,{\rm Im}\,\tau)^{-2}\sim\lambda\,,\theta_{3}\simeq\frac{|\alpha_{e}\alpha_{e2}Y^{(4)}_{\bf 1}|}{|\alpha_{\mu}|^{2}}(2\,{\rm Im}\,\tau)^{-2}\Delta^{26}_{\chi}\,.$ (134)
Here $\theta_{1}\sim{\cal O}(\lambda^{2})$, $\theta_{2}\sim{\cal O}(\lambda)$, and $\theta_{3}\ll{\cal O}(\lambda^{4})$ are roughly expected for $\Delta_{\chi}=0.634$ and ${\rm Im}\,\tau\simeq 1$. The right-handed charged-lepton mixing matrix $V_{R}^{\ell}$ is approximately the unit matrix up to ${\cal O}(\lambda^{4})$:
$\theta^{R}_{1}\simeq\frac{|\alpha_{\mu}\alpha_{\tau 2}Y^{(8)}_{\bf 1}|}{|\alpha_{\tau}|^{2}}(2\,{\rm Im}\,\tau)^{-4}\Delta^{7}_{\chi}\,$,
$\theta^{R}_{2}\simeq\frac{|\alpha_{e}\alpha_{\tau 1}Y^{(4)}_{\bf 1}|}{|\alpha_{\tau}|^{2}}(2\,{\rm Im}\,\tau)^{-2}\Delta^{20}_{\chi}\,$,
$\theta^{R}_{3}\simeq\frac{|\alpha_{e2}Y^{(4)}_{\bf 1}|}{|\alpha_{\mu}|}(2\,{\rm Im}\,\tau)^{-2}\Delta^{13}_{\chi}\,.$ (135)